You have obtained a gas chromatogram of a two component mixture of compounds A and B. The height of peak A is 3.74 cm. and the width at half height is 1.5 cm. The height of peak B is 3.6 cm. and the width at half height is 1.06 cm. The percentage of a component in the mixture is proportional to peak area. Estimate peak areas by multiplying height by width at half height. What was the percentage of compound B in the mixture? (Enter your answer as a number without any percent symbol or the word percent.)
Answer #1

The area of ​​each compound is calculated:

A = l * w = 3.74 * 1.5 = 5.61 cm2

B = 3.6 * 1.06 = 3.82 cm2

Total = 9.43 cm2

The percentage of B in the mixture is calculated:

% B = area B * 100 / total area = 3.82 * 100 / 9.43 = 40.5

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.